determine the equation of the graph of the given parabola

Diretrix = -3

focus= (3,-1)

vertex= (3,-2)
Answer #1

Given vertex (3,-2) focus (3,-1). Equation to find parabole is (x-h)² = 4P (Y-K). verted (3,-2 h => (x-3)? = 48 (Y-(-2)) (x²7(or) 2 Y = x 3X t 1 h 7